A speedboat starts from rest and accelerates at +2.20 m/s2 for 8.87 s. At the end of this time, the boat continues for an additional 4.43 s with an acceleration of +0.626 m/s2. Following this, the boat accelerates at -0.983 m/s2 for 6.32 s. (a) What is the velocity of the boat at t = 19.62 s? (b) Find the total displacement of the boat.
Answer #1

velocity after 8.87 s

v1 = 2.2* 8.87 = 19.514 m/s

displacement travelled

d1 = 0.5* 2.2* 8.87^2 = 86.545 m

velocity after another 4.43 s

v2 = 19.514 + 0.626* 4.43 = 22.287 m/s

displacement travelled

d2 = 19.514* 4.43 + 0.5* 0.626* 4.43^2

d2 = 92.59 m

velocity after further 6.32 s

v = 22.287 - 0.983* 6.32 = 16.07 ===>>> (Ans(a))

displacement travelled

d3 = 22. 287* 6.32 - 0.5* 0.983* 6.932^2

d3 = 117.235 m

net displacement

d = d1 + d2 + d3

d = 296.37 m

=====

comment before rate in case any doubt, will reply for sure.. goodluck